OpenStudy (anonymous):

Choose the correct simplification of (4x^3 + 3x^2 − 6x) − (10x^3 + 3x^2).

OpenStudy (kainui):

First off, you can distribute the negative sign to the whole second part in parenthesis, and from there you can add like terms, for example x^3+2x^3=3x^3 but you can't add x^3 to x^2

OpenStudy (anonymous):

1. 6x3 + 6x 2. −6x3 − 6x2 − 6x 3. −6x3 − 6x 4. 6x3 + 6x2 + 6x

OpenStudy (anonymous):

those are the multiple choice answers

OpenStudy (kainui):

Why can you only add x^2 to an x^2 and not an x with an x^2? Well try plugging in a number to see if they're equal. x^2+x^2=2x^2 and if you plug in a number, like 2, you get 4+4=8 which works. Now what about checking x^2+x does that equal x^3? x^2+x=x^3 what if you plug in a 2 now 4+2=8 since 6 and 8 aren't equal, you know you can't do that. =D